# Range, not Independence, Drives Modularity in Biological Inspired Representations

### v2

@@ 1 Introduction
Our brains are modular. At the macroscale, different regions, such as visual or language cortex, perform specialised roles; at the microscale, single neurons often precisely encode single variables such as self-position [17] or the orientation of a visual edge [24]. This mysterious alignment of meaningful concepts with single neuron activity has for decades fuelled hope for understanding a neuron’s function by finding its associated concept. Yet, as neural recording technology has improved, it has become clear that many neurons behave in ways that elude such simple categorisation: they appear to be mixed selective, responding to a mixture of variables in linear and nonlinear ways [39, 52]. The modules vs. mixtures debate has recently been reprised in the machine learning community. Both the disentangled representation learning and mechanistic interpretability subfields are interested in when neural network representations decompose into meaningful components. Findings have been similarly varied, with some studies showing meaningful single unit response properties and others clearly showing mixed tuning (for a full discussion on related work see Appendix A). This brings us to our main research question: Why are neurons, biological and artificial, sometimes modular and sometimes mixed selective?
In this work, we develop a theory that precisely determines, for any dataset, whether the optimal learned representations will be modular or not. More precisely, in the linear autoencoder setting, we show that modularity in biologically constrained representations is governed by a sufficient spread condition that can roughly be thought of as measuring the extent to which the support of the source variables (sources, aka factors of variation) is rectangular. This sufficient spread condition bears resemblance to identifiability conditions derived in the matrix factorisation literature [49, 50], though both the precise form of the problem we study and the condition we derive differ (Appendix A). This condition on the sources is much weaker than the case of mutual independence studied in previous work [56], and commensurately broadens the settings we can understand using our theory. For example, the loosening from statistical independence to rectangular support enables us to predict when linear recurrent neural network (RNN) representations of dynamic variables modularise.
@@ 2.1 Preliminaries
Let ${\bm{s}}\in\mathbb{R}^{d_{s}}$ be a vector of $d_{s}$ scalar source variables (sources, aka factors of variation). We are interested in how the empirical distribution, $p({\bm{s}})$, affects whether the sources’ neural encoding (latents), ${\bm{z}}\in\mathbb{R}^{d_{z}}$, are modular with respect to the sources, i.e., whether each neuron (latent) is functionally dependent on at most one source. Following [56], we build a simplified model in which neural firing rates perfectly linearly autoencode the sources while maintaining nonnegativity (since firing rates are nonnegative):
@@
Subject to the above constraints, we study the representation that uses the least energy, as in the classic efficient coding hypothesis [3]. We quantify this using the $\ell^{2}$ norm of the firing rates and weights (other activity norms considered later):
@@
We remark that there are common analogues of representational nonnegativity and weight energy efficiency in modern machine learning (ReLU activation functions and weight decay, respectively). When the sources are statistically independent, i.e. $p({\bm{s}})=\prod_{i=1}^{d_{s}}p(s_{i})$, then the optima of Equation 2 have modular latents [56]. We now improve on this result by showing necessary and sufficient conditions that guarantee modularisation for any dataset, not just those that have statistically independent sources.
@@ 2.2 Intuition for Source Support Modularisation Conditions
that is functionally dependent on both $s_{1}$ and $s_{2}$. This neuron is mixed selective. Perhaps, however, a modular representation, in which this neuron is broken into two separate modular encodings, would be better. If so, the optimal solution would have two neurons each coding for a single source:
@@
where the minimizations are constrained to the support of the empirical distribution $p(s_{1},s_{2})$. Now that both representations are specified, we can compare their costs. When calculating the activity energy one finds that, in this simplified setting, most terms are zero or cancel, and hence the mixed selective  Equation 3 case uses less energy than the modular  Equation 4 when
@@
In this case, the energy of the mixed solution will always be worse than the modular, since $b_{j}^{2}=(b_{j^{\prime}}+b_{j^{\prime\prime}})^{2}>b_{j^{\prime}}^{2}+b_{j^{\prime\prime}}^{2}$. Alternatively, mixing will be preferred when $s_{1}$ and $s_{2}$ do not take on their most negative values at the same time, as in the bottom row of  Figure 1a, since then $b_{j}$ does not have to be as large to maintain positivity, and the corresponding energy saving satisfies the key inequality Equation 6.
Figure 1: a) Top row: Values of two sources across a dataset and their modular encoding with associated costs (shaded regions). Middle row: If the dataset includes the red datapoint (left) then the two variables take their minimal value at the same time, and a mixed encoding must include a large bias, using more energy (right). Bottom row: If the red point is instead missing (left) then the mixed encoding can use a smaller bias while remaining positive and save energy (right). b) Our modularisation conditions (Theorem 2.1) are equivalent to whether the convex hull of the data (orange regions) encloses a data-derived ellipse (Theorem 2.2). The precision of these conditions is validated in two experiments (rows) in which the source support differs by a single datapoint. Removing this datapoint stops the convex hull from enclosing the key ellipse, resulting in a mixed-selective neuron. In the middle column, we plot the weight vectors of the neurons. The top representation is modular as all weight vectors are axis aligned, whereas in the bottom there are mixed selective neurons. This is reiterated by extracting the most mixed selective neuron (orange dot) and plotting its response in the right column. c) Our conditions (left) predict modularisation better than measures of statistical independence such as source multiinformation (right). Here the y axis measures how mixed the representation is by the largest angle between a neuron’s weight vector and one of the source axes.
@@ 2.3 Precise Conditions for Modularising Biologically Inspired Linear Autoencoders
Our theory prescribes a set of inequalities (Equation 9) that determine whether an optimal representation is modular. These inequalities come from the difference in activity energy between a modular and mixed solutions, just like the intuition we built up in Section 2.2, and in particular Equation 6. If a single inequality is broken, the representation is mixed (at least in part); else, the representation is modular: each neuron’s activity is a function of a single source. These inequalities depend on two key properties of the sources: the shape of the source distribution support in extremal regions, and the pairwise source correlations. Remarkably, they do not depend on the energy tradeoff parameter $\lambda$. These inequalities can be visualised, as done in the wrapped figure below. For a given dataset, $\{{\bm{s}}^{[i]}\}$ (blue dots), we can calculate $\left\langle\bar{s}^{[i]}_{j}\bar{s}^{[i]}_{j^{\prime}}\right\rangle_{i}$ and each $\min_{i}\bar{s}^{[i]}_{j}$ as they are simple functions of the dataset. For all unit ${\bm{w}}$, we can draw the line
${\bm{w}}^{\top}{\bm{x}}=\sqrt{\sum_{j=1}^{d_{s}}\left(w_{j}\min_{i}\bar{s}^{[i]}_{j}\right)^{2}-\sum^{d_{s}}_{j,j^{\prime}\neq j}w_{j}w_{j^{\prime}}\left\langle\bar{s}^{[i]}_{j}\bar{s}^{[i]}_{j^{\prime}}\right\rangle_{i}}=\sqrt{{\bm{w}}^{T}{\bm{S}}{\bm{w}}},$ (10)
@@
${\bm{S}}=\begin{bmatrix}(\min_{i}s_{1}^{[i]})^{2}&-\langle s_{1}^{[i]}s_{2}^{[i]}\rangle_{i}&-\langle s_{1}^{[i]}s_{3}^{[i]}\rangle_{i}&\ldots\\ -\langle s_{1}^{[i]}s_{2}^{[i]}\rangle_{i}&(\min_{i}s_{2}^{[i]})^{2}&-\langle s_{2}^{[i]}s_{3}^{[i]}\rangle_{i}&\ldots\\ -\langle s_{1}^{[i]}s_{3}^{[i]}\rangle_{i}&-\langle s_{2}^{[i]}s_{3}^{[i]}\rangle_{i}&(\min_{i}s_{3}^{[i]})^{2}&\ldots\\ \vdots&\vdots&\vdots&\ddots\end{bmatrix}$ (11)
Each ${\bm{w}}$ gives us a line, and if there exists a line (such as the red one) that bounds the source support, then an optimal representation must mix. This exercise also motivates the following equivalent statement of our conditions.
@@
In the same setting as Theorem 2.1, define the set $E=\{{\bm{y}}:{\bm{y}}^{T}{\bm{S}}^{-1}{\bm{y}}=1\}$. Then an equivalent statement of Theorem 2.1 is the representation modularises iff $E$ lies within the convex hull of the datapoints.
This therefore provides a simple test: create ${\bm{S}}$ and draw the set $E$ which, if ${\bm{S}}$ is positive definite (as it often is), is an ellipse as shown in the wrapped figure above. The optimal representation modularises iff the convex hull of the datapoints encloses $E$, figure Figure 1b. If ${\bm{S}}$ has some positive and some negative eigenvalues, then the optimal representation must mix.
The above results apply to autoencoders trained to linearly encode and decode the pure source vector ${\bm{s}}^{[i]}$. In Section B.4 we derive similar results for linear mixtures of sources, i.e. where the data the autoencoder receives and reconstructs is ${\bm{x}}^{[i]}={\bm{A}}{\bm{s}}^{[i]}$ for some mixing matrix ${\bm{A}}$.
@@ 2.4 Range Independent Variables Modularise
As presented, our theory has some limitations. It focuses on sources that are 1-dimensional, and uses a specific choice of activity norm, the L2, when others might be more reasonable. In fact, however, the core result that it is the independence of the variables’ support (i.e. how rectangular they are) that drives modularity is very broadly true. In Appendix C we show that this core result generalises to other activity norms, and in Appendix D we show that it also applies to the modularisation of multi-dimensional variables such as angles on a 2D circle. Thus, our key takeaway is that the independence of support is what determines whether the optimal representation is modular.
@@ 2.5 Empirical Tests
Validation of linear autoencoder theory. We show our inequalities correctly predict modularisation. In particular, as an illustration of the precision of our theory, we create a dataset which transitions from inducing modularising to inducing mixing via the removal of a single critical datapoint (Figure 1b). Further, we generate many datasets, create the optimal biological representation, and measure the angle $\theta$ between the most-mixed neuron’s weight vector and its closest source direction, a proxy for modularity. The left of Figure 1c shows that our theory correctly predicts which datasets are modular ($\theta=0$). Further, despite our theory being binary (will it modularise or not?), empirically we see that a good proxy for how mixed the optimal representation will be is the degree to which the inequalities in Theorem 2.1 are broken. Finally, on the right of Figure 1c we show that on the same datasets a measure of source statistical interdependence, as used in previous work, does not predict modularisation.
Predictions beyond our theory. From our theory we extract qualitative trends to empirically test in more complex settings. Since extremal points play an outsized role in determining modularisation, we consider three trends that highlight these effects. (1) Datasets from which successively smaller corners have been removed should become successively less mixed, until at a critical threshold the representation modularises. (2) It is vital that not just any data, but rather corner slices, that are removed. Removing similar amounts of random or centrally located data from the dataset should not cause as much mixing. (3) Introducing correlations into a dataset while preserving extreme-point or range independence should preserve modularity relatively well.
@@ 3 Modularisation in Biologically Inspired Nonlinear Feedforward Networks
Figure 2: a) We study three tasks of varying source dimensionality $d_{s}$, model complexity, supervision signal, and input data modality (Section 3). For each task, we plot conditional InfoM, a measure of modularity, against normalised source multiinformation (NSI), a measure of the statistical dependencies in the sources. Across the three tasks, we observe that cutting corners severely degrades modularity (orange), whereas equivalently changing the NSI in ways that preserve the rectangular support boundary (green, purple) affects modularity less. Shaded regions denote standard error of the mean. b) For What-Where Regression, we visualise the source distributions at various NSI values (top) as well as the tuning of an example neuron (bottom; vertical: What; horizontal: Where).
@@
Metrics for representational modularity and inter-source statistical dependence. To quantify the modularity of a nonlinear representation, we design a family of metrics called conditional information-theoretic modularity (CInfoM), an extension of the InfoM metric proposed by Hsu et al. [22]. Intuitively, a representation is modular if each neuron is informative about only a single source. We therefore calculate the conditional mutual information between a neuron’s activity and each source variable given all other sources. The conditioning is necessary to remove the effect of sources leaking information about each other; prior works consider independent sources or do not account for this effect. CInfoM then measures the extent to which a neuron specialises to its favourite source, relative to its informativeness about all sources. In order to compare multiple schemes that change $p({\bm{s}})$ in different ways, we report normalised source multiinformation (NSI) as a measure of source statistical interdependence. NSI involves estimating the source multiinformation (aka total correlation) $D_{\text{KL}}(p({\bm{s}})\;\|\;\textstyle\prod_{i=1}^{d_{s}}p(s_{i}))$ and normalizing by $\sum_{i=1}^{d_{s}}H(s_{i})-\max_{i}H(s_{i})$. We defer further exposition of the CInfoM metric to Appendix F.
@@
What-where regression. Inspired by the modularisation of what and where into the ventral and dorsal visual streams, we present nonlinear networks with simple binary images in which one pixel is on. The network is trained to concurrently report within which region of the image (“where”), and where within that region (“what”) the on pixel is found, producing two outputs which we take as our sources, each an integer between one and nine. (More complex inputs or one-hot labels also work, see Appendix G.) We regularise the activity and weight energies, and enforce nonnegativity using a ReLU. If what and where are independent from one another, e.g. both uniformly distributed, then under our biological constraints (but not without them, see Appendix G) the hidden layer modularises what from where. Breaking the independence of what and where leads to mixed representations in patterns that qualitatively agree with our theory (Figure 2a left): cutting corners from the source support causes increasing mixing. Conversely, making other more drastic changes to the support, such as removing the diagonal, does not induce mixing. Similarly, introducing source correlations while preserving the rectangular support introduces less mixing when compared to corner cutting that induces the same amount of mutual information between what and where. The various source distributions at different NSI values are visualised in Figure 2b.
@@
Nonlinear autoencoding of sources. Next we study a nonlinear autoencoder trained to autoencode multidimensional source variables. Again we find (Figure 2a middle) that under biological constraints independent source variables ($\text{NSI}=0$) result in modular latents and that source corner cutting (orange) induces far more latent mixing compared to introducing source correlations while preserving rectangular support (green), or removing central data (purple).
@@
Disentangled representation learning of images. Finally, for an experiment involving high-dimensional image data, we turn to a recently introduced state-of-the-art disentangling method, quantised latent autoencoding (QLAE; Hsu et al. [22, 23]). QLAE is the natural machine learning analogue to our biological constraints. It has two components: (1) the weights in QLAE are heavily regularised, like our weight energy loss, and (2) the latent space is axis-wise quantised, introducing a privileged basis with low channel capacity. In our biologically inspired networks, nonnegativity and activity regularisation conspire to similarly structure the representation: nonnegativity creates a preferred basis, and activity regularisation encourages the representation to use as small a portion of the space as possible. We study the performance of QLAE trained to autoencode a subset of the Isaac3D dataset [36], a naturalistic image dataset with well defined underlying latent dimensions. We find the same qualitative patterns: corner cutting is a more important determinant of mixing than range-independent correlations or the removal of centrally located data (Figure 2a right).
4 Modularisation in Biologically Inspired Recurrent Neural Networks

### v4

# Range, not Independence, Drives Modularity in Biologically Inspired Representations

@@ 1 Introduction
Our brains are modular. At the macroscale, different regions, such as visual or language cortex, perform specialised roles; at the microscale, single neurons often precisely encode single variables such as self-position (Hafting et al., 2005) or the orientation of a visual edge (Hubel and Wiesel, 1962). This mysterious alignment of meaningful concepts with single neuron activity has for decades fuelled hope for understanding a neuron’s function by finding its associated concept. Yet, as neural recording technology has improved, it has become clear that many neurons behave in ways that elude such simple categorisation: they appear to be mixed selective, responding to a mixture of variables in linear and nonlinear ways (Rigotti et al., 2013; Tye et al., 2024). The modules vs. mixtures debate has recently been reprised in the machine learning community. Both the disentangled representation learning and mechanistic interpretability subfields are interested in when neural network representations decompose into meaningful components. Findings have been similarly varied, with some studies showing meaningful single unit response properties and others clearly showing mixed tuning (for a full discussion on related work see Appendix A). This brings us to our main research question: Why are neurons, biological and artificial, sometimes modular and sometimes mixed selective?
In this work, we develop a theory that precisely determines, for any dataset, whether the optimal learned representations will be modular or not. More precisely, in the linear autoencoder setting, we show that modularity in biologically constrained representations is governed by a sufficient spread condition that can roughly be thought of as measuring the extent to which the range of the source variables (sources, aka factors of variation) is rectangular. This sufficient spread condition bears resemblance to identifiability conditions derived in the matrix factorisation literature (Tatli and Erdogan, 2021a; b), though both the precise form of the problem we study and the condition we derive differ (Appendix A). This condition on the sources is much weaker than the case of mutual independence studied in previous work (Whittington et al., 2023), and commensurately broadens the settings we can understand using our theory. For example, the loosening from statistical independence to rectangular support enables us to predict when linear recurrent neural network (RNN) representations of dynamic variables modularise.
@@ 2.1 Preliminaries
Let ${\bm{s}}\in\mathbb{R}^{d_{s}}$ be a vector of $d_{s}$ scalar source variables (sources, aka factors of variation). We are interested in how the empirical distribution, $p({\bm{s}})$, affects whether the sources’ neural encoding (latents), ${\bm{z}}\in\mathbb{R}^{d_{z}}$, are modular with respect to the sources, i.e., whether each neuron (latent) is functionally dependent on at most one source. Following Whittington et al. (2023), we build a simplified model in which neural firing rates perfectly linearly autoencode the sources while maintaining nonnegativity (since firing rates are nonnegative):
@@
Subject to the above constraints, we study the representation that uses the least energy, as in the classic efficient coding hypothesis (Barlow et al., 1961). We quantify this using the $\ell^{2}$ norm of the firing rates and weights (other activity norms considered later):
@@
We remark that there are common analogues of representational nonnegativity and weight energy efficiency in modern machine learning (ReLU activation functions and weight decay, respectively). When the sources are statistically independent, i.e. $p({\bm{s}})=\prod_{i=1}^{d_{s}}p(s_{i})$, then the optima of Equation 2 have modular latents (Whittington et al., 2023). We now improve on this result by showing necessary and sufficient conditions that guarantee modularisation for any dataset, not just those that have statistically independent sources.
@@ 2.2 Intuition for Source Support Modularisation Conditions
It is functionally dependent on both $s_{1}$ and $s_{2}$, i.e. it is mixed selective. Perhaps, however, a modular representation, in which this neuron is broken into two separate modular encodings, would be better. We can create such a solution with two neurons each coding a single source:
@@
where the minimisations are over the empirical distribution $p(s_{1},s_{2})$. Now that both representations are specified, we can compare their costs. In our problem modularity is driven only by the activity loss (Appendix B). Further, in this simplified setting, most terms in the activity loss are zero or cancel, and one finds that the mixed selective Equation 3 case uses less energy than the modular Equation 4 when
@@
In this case, the energy of the mixed solution will always be worse than the modular, since $b_{j}^{2}=(b_{j^{\prime}}+b_{j^{\prime\prime}})^{2}>b_{j^{\prime}}^{2}+b_{j^{\prime\prime}}^{2}$. Alternatively, mixing will be preferred when $s_{1}$ and $s_{2}$ do not take on their most negative values at the same time, as in the bottom row of  Figure 1a, since then $b_{j}$ can be smaller while maintaining positivity, and the corresponding energy saving satisfies the key inequality Equation 6.
Figure 1: a) Top row: Values of two sources across a dataset and their modular encoding with associated costs (shaded regions). Middle row: If the dataset includes the red datapoint (left) then the two variables take their minimal value at the same time, and the mixed encoding must include a large bias, using more energy (right). Bottom row: If the red point is instead missing (left) then the mixed encoding can use a smaller bias while remaining positive and save energy (right). b) Our modularisation conditions (Theorem 2.1) are equivalent to whether the convex hull of the data (orange regions) encloses a data-derived ellipse (Theorem 2.2). The precision of these conditions is validated in two experiments (rows) in which the source support differs by a single datapoint. Removing this datapoint stops the convex hull from enclosing the key ellipse, resulting in a mixed-selective representation. In the middle column, we plot the weight vectors of the neurons. The top representation is modular as all weight vectors are axis aligned, whereas the bottom contains mixed selective neurons. This is reiterated by extracting the most mixed selective neuron (orange dot) and plotting its activity as a function of the sources (right column). c) Our conditions (left) predict modularisation better than measures of statistical independence such as source multiinformation (right). Here the y axis measures how mixed the representation is by the largest angle between a neuron’s weight vector and one of the source axes, and the left x axis measures how much the inequalities are broken by - a rough heuristic for how mixed the optimal solution is.
@@ 2.3 Precise Conditions for Modularising Biologically Inspired Linear Autoencoders
Our theory prescribes a set of inequalities Equation 9 that determine whether an optimal representation is modular. These inequalities come from the difference in activity energy between a modular and mixed solutions, just like the intuition we built up in Section 2.2, and in particular Equation 6. If a single inequality is broken, the optimal representation is mixed (at least in part); else, it is modular: optimally, each neuron’s activity is a function of a single source. These inequalities depend on two key properties of the sources: the shape of the source distribution’s support in extremal regions, and the pairwise source correlations. Remarkably, they do not depend on the energy tradeoff parameter $\lambda$. These inequalities can be visualised, as done in the wrapped figure below. For a given dataset, $\{{\bm{s}}^{[i]}\}$ (blue dots), we can calculate $\left\langle\bar{s}^{[i]}_{j}\bar{s}^{[i]}_{j^{\prime}}\right\rangle_{i}$ and each $\min_{i}\bar{s}^{[i]}_{j}$ as they are simple functions of the dataset. For all unit ${\bm{w}}$, we can draw the line
${\bm{w}}^{\top}{\bm{x}}=\sqrt{\sum_{j=1}^{d_{s}}\left(w_{j}\min_{i}\bar{s}^{[i]}_{j}\right)^{2}-\sum^{d_{s}}_{j,j^{\prime}\neq j}w_{j}w_{j^{\prime}}\left\langle\bar{s}^{[i]}_{j}\bar{s}^{[i]}_{j^{\prime}}\right\rangle_{i}}=\sqrt{{\bm{w}}^{T}{\bm{F}}{\bm{w}}},$ (10)
@@
${\bm{F}}=\begin{bmatrix}(\min_{i}s_{1}^{[i]})^{2}&-\langle s_{1}^{[i]}s_{2}^{[i]}\rangle_{i}&-\langle s_{1}^{[i]}s_{3}^{[i]}\rangle_{i}&\ldots\\ -\langle s_{1}^{[i]}s_{2}^{[i]}\rangle_{i}&(\min_{i}s_{2}^{[i]})^{2}&-\langle s_{2}^{[i]}s_{3}^{[i]}\rangle_{i}&\ldots\\ -\langle s_{1}^{[i]}s_{3}^{[i]}\rangle_{i}&-\langle s_{2}^{[i]}s_{3}^{[i]}\rangle_{i}&(\min_{i}s_{3}^{[i]})^{2}&\ldots\\ \vdots&\vdots&\vdots&\ddots\end{bmatrix}$ (11)
Each ${\bm{w}}$ gives us a line, and if there is at least one line that bounds the source support (such as the red one), then an inequality is broken and the optimal representation is mixed. Else it will be modular. This exercise also motivates the following equivalent statement of our conditions.
@@
In the same setting as Theorem 2.1, define the set $E=\{{\bm{y}}:{\bm{y}}^{T}{\bm{F}}^{-1}{\bm{y}}=1\}$. Then an equivalent statement of Theorem 2.1 is the representation modularises iff $E$ lies within the convex hull of the datapoints.
This therefore provides a simple test: create ${\bm{F}}$ and draw the set $E$ which, if ${\bm{F}}$ is positive definite (as it often is), is an ellipse as shown in the wrapped figure above. The optimal representation modularises iff the convex hull of the datapoints encloses $E$Figure 1b. If ${\bm{F}}$ has some positive and some negative eigenvalues, then the set $E$ is unbounded, and the optimal representation must mix.
@@ 2.4 Range Independent Variables Modularise
As presented, our theory has some limitations. It focuses on sources that are 1-dimensional, and uses a specific choice of activity norm, the L2, when others might be more reasonable. In fact, however, the core result that it is the independence of the variables’ range (i.e. how rectangular they are) that drives modularity is very broadly true. In Section B.4 we show that this core result generalises to other activity norms, and in Section B.5 we show that it also applies to the modularisation of multi-dimensional variables such as angles on a 2D circle. Thus, our key takeaway is that the independence of variables’ range is what determines whether the optimal representation is modular.
@@ 2.5 Empirical Tests
Validation of linear autoencoder theory. We show our inequalities correctly predict modularisation. In particular, as an illustration of the precision of our theory, we create a dataset which transitions from inducing modularising to mixing via the removal of a single critical datapoint (Figure 1b). Further, we generate many datasets, create the optimal representation, and measure the angle $\theta$ between the most-mixed neuron’s weight vector and its closest source direction, a proxy for modularity. The left of Figure 1c shows that our theory correctly predicts which datasets are modular ($\theta=0$). Further, despite our theory being binary (will it modularise or not?), empirically we see that the degree to which the inequalities in Theorem 2.1 are broken is a good proxy for how mixed the optimal representation is. Finally, on the right of Figure 1c we show that on the same datasets a measure of source statistical interdependence, as used in previous work, does not predict modularisation.
Predictions beyond our theory. From our theory we extract qualitative trends to empirically test in more complex settings. Since extremal points play an outsized role in determining modularisation, we consider three trends that highlight these effects. (1) Datasets from which successively smaller corners have been removed should become successively less mixed, until at a critical threshold the representation modularises. (2) It is vital that it is not just any data but corner slices that are removed. Removing similar amounts of random or centrally located data from the dataset should not cause as much mixing. (3) Introducing correlations into a dataset while preserving extreme-point or range independence should preserve modularity relatively well.
@@ 3 Modularisation in Biologically Inspired Nonlinear Feedforward Networks
Figure 2: a) We study three tasks of varying source dimensionality $d_{s}$, model complexity, supervision signal, and input data modality (Section 3). For each task, we plot conditional InfoM, a measure of modularity, against normalised source multiinformation (NSI), a measure of the statistical dependencies in the sources. Across the three tasks, we observe that cutting corners severely degrades modularity (orange), whereas equivalently changing the NSI in ways that preserve the rectangular support boundary (green, purple) affects modularity less. Shaded regions denote standard error of the mean. b) For What-Where Regression, we visualise the source distributions at various NSI values (top) as well as the tuning of an example neuron. The $x$ ($y$) axes correspond to different what (where) values, and the colour shows the neuron’s response. All neurons begin modular (left three plots), dropping corners heavily degrades modularity (left plot of subsequent groups of three).
@@
Metrics for representational modularity and inter-source statistical dependence. To quantify the modularity of a nonlinear representation, we design a family of metrics called conditional information-theoretic modularity (CInfoM), an extension of the InfoM metric proposed by Hsu et al. (2023). Intuitively, a representation is modular if each neuron is informative about only a single source. We therefore calculate the conditional mutual information between a neuron’s activity and each source variable given all other sources. The conditioning on other sources is necessary to remove the effect of sources leaking information about each other; prior works consider independent sources or do not account for this effect. CInfoM then measures the extent to which a neuron specialises to its favourite source, relative to its informativeness about all sources. In order to compare multiple schemes that change $p({\bm{s}})$ in different ways, we report normalised source multiinformation (NSI) as a measure of source statistical interdependence. NSI involves estimating the source multiinformation (aka total correlation) $D_{\text{KL}}(p({\bm{s}})\;\|\;\textstyle\prod_{i=1}^{d_{s}}p(s_{i}))$ and normalizing by $\sum_{i=1}^{d_{s}}H(s_{i})-\max_{i}H(s_{i})$. We defer further exposition of the CInfoM metric to Appendix E.
@@
What-where regression. Inspired by the modularisation of what and where into the ventral and dorsal visual streams, we present nonlinear networks with simple binary images in which one pixel is on. The network is trained to concurrently report within which region of the image (“where”), and where within that region (“what”) the on pixel is found, producing two outputs which we take as our sources, each an integer between one and nine. (More complex inputs or one-hot labels also work, see Appendix F.) We regularise the activity and weight energies, and enforce nonnegativity using a ReLU. If what and where are independent from one another, e.g. both uniformly distributed, then under our biological constraints (but not without them, see Appendix F) the hidden layer modularises what from where. Breaking the independence of what and where leads to mixed representations in patterns that qualitatively agree with our theory (Figure 2a left): cutting corners from the source support causes increasing mixing. Conversely, making other more drastic changes to the support, such as removing the diagonal, does not induce mixing. Similarly, introducing source correlations while preserving the rectangular support introduces less mixing when compared to corner cutting that induces the same amount of mutual information between what and where. The various source distributions at different NSI values are visualised in Figure 2b.
@@
Nonlinear autoencoding of sources. Next we study a nonlinear autoencoder trained to autoencode a set of source variables. Again we find (Figure 2a middle) that under biological constraints independent source variables ($\text{NSI}=0$) result in modular latents and that source corner cutting (orange) induces far more latent mixing compared to introducing source correlations while preserving rectangular support (green), or removing central data (purple).
@@
Disentangled representation learning of images. Finally, for an experiment involving high-dimensional image data, we turn to a recently introduced state-of-the-art disentangling method, quantised latent autoencoding (QLAE; Hsu et al. (2023; 2024)). QLAE is the natural machine learning analogue to our biological constraints. It has two components: (1) the weights in QLAE are heavily regularised, like our weight energy loss, and (2) the latent space is axis-wise quantised, introducing a privileged basis with low channel capacity. In our biologically inspired networks, nonnegativity and activity regularisation conspire to similarly structure the representation: nonnegativity creates a preferred basis, and activity regularisation encourages the representation to use as small a portion of the space as possible. We study the performance of QLAE trained to autoencode a subset of the Isaac3D dataset (Nie, 2019), a naturalistic image dataset with well defined underlying latent dimensions. We find the same qualitative patterns: corner cutting is a more important determinant of mixing than range-independent correlations or the removal of centrally located data (Figure 2a right).
4 Modularisation in Biologically Inspired Recurrent Networks